# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v1

@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
$\left\{\begin{matrix}\mathcal{D}_{2}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}\\ \|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G^{\prime}_{n},G_{*})\to 0,\end{matrix}\right.$
which indicates that $\|f_{G^{\prime}_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}\to 0$ as $n\to\infty$.
Note that $\Theta$ is a compact set, there exists a mixing measure $G^{\prime}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that one of $G^{\prime}_{n}$’s subsequences converges to $G^{\prime}$. Since $\mathcal{D}_{2}(G^{\prime}_{n},G_{*})>\varepsilon^{\prime}$, we deduce that $\mathcal{D}_{2}(G^{\prime},G_{*})>\varepsilon^{\prime}$. By invoking the Fatou’s lemma, we have that
Thus, we have $f_{G^{\prime}}=f_{G_{*}}$ for $\mu-$almost surely ${\bm{X}}$. From the identifiability property, we deduce that $G^{\prime}\equiv G_{*}$. It follows that $\mathcal{D}_{2}(G^{\prime},G_{*})=0$, contradicting the fact that $\mathcal{D}_{2}(G^{\prime},G_{*})>\varepsilon^{\prime}>0$. Hence, the proof is completed.
Identifiability property.
@@
We now show that if $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, then it follows that $G\equiv G_{*}$.
For any $G\in\mathcal{G}_{L^{\prime}}(\Xi)$ and $G_{*}$, let us denote
$\mathrm{softmax}_{G}(u)$ $=\dfrac{\exp(u)}{\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L^{\prime}}\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})},$
$\mathrm{softmax}_{G_{*}}(u_{*})$ $=\dfrac{\exp(u_{*})}{\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})},$
where
$u$$\in\{{\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j};(BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}}:j\in[N],j^{\prime}\in[L^{\prime}]\},$
$u_{*}$$\in\{{\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j};(BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}}:j\in[N],j^{\prime}\in[L]\}.$
Since $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, we have
$\sum_{j=1}^{N}\mathrm{softmax}_{G}({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L^{\prime}}\mathrm{softmax}_{G}((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})CW_{2}W_{1j^{\prime}}{\bm{X}}$
$=\sum_{j=1}^{N}\mathrm{softmax}_{G_{*}}({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L}\mathrm{softmax}_{G_{*}}((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})CW_{*,2}W_{*,1j^{\prime}}{\bm{X}}.$ (38)
Thus, we must have that $L=L^{\prime}$. As a result, we obtain that
$\{\mathrm{softmax}_{G}((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}}):j^{\prime}\in[L]\}$ =$=\{\mathrm{softmax}_{G_{*}}((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}}):j^{\prime}\in[L^{\prime}]\},$
for almost every ${\bm{X}}$. We may assume that
$\mathrm{softmax}_{G}((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})=\mathrm{softmax}_{G_{*}}((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}}),$
for any $j^{\prime}\in[L]$, for almost every ${\bm{X}}$. Since the softmax function is invariant to translation, this result indicates that $b_{j^{\prime}}=b_{*,j^{\prime}}+r$ for some $r\in\mathbb{R}$ and for any $j^{\prime}\in[L]$. Then, the equation (38) can be reduced to
@@
for almost every ${\bm{X}}$. Next, we will partition the index set $[L]$ into $m$ subsets $K_{1},K_{2},\ldots,K_{m}$ where $m\leq L$, such that $\exp{(b_{j})}=\exp{(b_{*,j^{\prime}})}$ for any $j,j^{\prime}\in K_{i}$ and $i\in[m]$. It follows that $\exp{(b_{j})}\neq\exp{(b_{*,j^{\prime}})}$ when $j,j^{\prime}$ do not belong to the same set $K_{i}$. Thus, we can rewrite equation (39) as
$\sum_{i=1}^{m}\sum_{j\in K_{i}}\exp{(b_{j})}\exp{((BW_{2}\sigma(W_{1j}{\bm{X}}))^{\top}{\bm{X}})}CW_{2}W_{1j}{\bm{X}}$
$=\sum_{i=1}^{m}\sum_{j^{\prime}\in K_{i}}\exp{(b_{*,j})}\exp{((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}})}CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$
for almost surely ${\bm{X}}$. Given the above equation, for each $i\in[m]$, we obtain that
$\{W_{2}W_{1j}{\bm{X}}:j\in K_{i}\}=\{W_{*,2}W_{*,1j^{\prime}}{\bm{X}}:j^{\prime}\in K_{i}\},$
Consequently, we get that
$\sum_{i=1}^{m}\sum_{j\in K_{i}}\exp{(b_{j})}\delta_{W_{2}W_{1j}}=\sum_{i=1}^{m}\sum_{j^{\prime}\in K_{i}}\exp{(b_{*,j})}\delta_{W_{*,2}W_{*,1j^{\prime}}},$
which implies that $G\equiv G_{*}$. As a consequence, the proof is completed. ∎
@@ Appendix C Related Work
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, [23] focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, [48] established convergence rates for both density and parameter estimation in Softmax gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, [46, 47] employed least squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.

### v3

@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
$\left\{\begin{matrix}\mathcal{D}_{2}(G^{\prime}_{n},\widetilde{G}_{*})>\varepsilon^{\prime}\\ \|f_{G^{\prime}_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G^{\prime}_{n},\widetilde{G}_{*})\to 0,\end{matrix}\right.$
as long as $n$ approaches the infinity. These results prove that $\|f_{G^{\prime}_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}\to 0$ as $n$ goes to infinity.
From the hypothesis, the parameter space $\Theta$ is compact. Therefore, one of $G^{\prime}_{n}$’s subsequences converges to some mixing measure $G^{\prime}$ where $G^{\prime}$ lies in the space $\mathcal{G}_{L^{\prime}}(\Theta)$. From the hypothesis, we have $\mathcal{D}_{2}(G^{\prime}_{n},\widetilde{G}_{*})>\varepsilon^{\prime}$. By taking the limit of both sides as $n\to\infty$, we obtain that $\mathcal{D}_{2}(G^{\prime},G_{*})\geq\varepsilon^{\prime}$.
An application of the Fatou’s lemma leads to the following result:
$0=\lim_{n\to\infty}\|f_{G^{\prime}_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}\geq\int\liminf_{n\to\infty}\left\|f_{G^{\prime}_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right\|^{2}d\mu({\bm{X}}).$
This inequality is only possible if $f_{G^{\prime}}=f_{\widetilde{G}_{*}}$ for almost surely ${\bm{X}}$.
According to the identifiability of the function $f_{G}({\bm{X}})$, that equation only holds when $G^{\prime}\equiv\widetilde{G}_{*}$. As a consequence, we obtain that $\mathcal{D}_{2}(G^{\prime},\widetilde{G}_{*})=0$. It contradicts to the assumption that $\mathcal{D}_{1}(G^{\prime},\widetilde{G}_{*})\geq\varepsilon^{\prime}>0$. Hence, the proof of the global part is completed. We achieve the conclusion of the theorem.
Proof for the identifiability property.
@@
The key claim that we aim to show is that if the equation $f_{G}({\bm{X}})=f_{\widetilde{G}_{*}}({\bm{X}})$ for almost every ${\bm{X}}$, then we achieve that $G\equiv\widetilde{G}_{*}$, namely, the two mixing measures are identical.
From the hypothesis, as $f_{G}({\bm{X}})=f_{G_{*}}({\bm{X}})$ for almost all ${\bm{X}}$, we achieve that
$\sum_{j=1}^{N}\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))}{\tilde{D}_{f,G}({\bm{X}})}h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{\tilde{L}}\frac{\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})}{\tilde{D}_{f,G}({\bm{X}})}CW_{2}W_{1j^{\prime}}{\bm{X}}$
$=\sum_{j=1}^{N}\frac{\exp({\bm{X}}^{\top}A^{0}_{j}{\bm{X}}+a^{0}_{j}))}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})}h({\bm{X}},\eta^{0}_{j})+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})}CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$ (39)
where $G=\sum_{j=1}^{\tilde{L}}\exp(b_{j})\delta_{W_{2}W_{1j}}$. Furthermore, we define
$\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})$ $=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{L}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}}),$
$\tilde{D}_{f,G}({\bm{X}})$ $=\sum_{k=1}^{N}\exp({\bm{X}}^{\top}A^{0}_{k}{\bm{X}}+a^{0}_{k})+\sum_{j^{\prime}=1}^{\tilde{L}}\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}}).$
That equation implies that the number of atoms of $G$ and $\widetilde{G}_{*}$ should be identical, namely, we have $L=\tilde{L}$. Therefore, the following result holds
$\Big{\{}\frac{\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})}{\tilde{D}_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Big{\}}$ =$=\Big{\{}\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})}:j^{\prime}\in[L]\Big{\}},$
for almost surely ${\bm{X}}$. By relabelling the indices of these two sets, we can assume without loss of generality that
$\frac{\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})}{\tilde{D}_{f,G}({\bm{X}})}$ $=\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})},$
for any index $j^{\prime}\in[L]$ and for almost surely ${\bm{X}}$. From the invariance to translation property of the softmax function, the equation (39) becomes
@@
for almost surely ${\bm{X}}$. This equation suggests that there exists a partition $K_{1},K_{2},\ldots,K_{m}$ of the set $[L]$ for some $m$ such that we have $\exp(b_{j_{1}})=\exp(b_{*,j_{2}})$ for any $j_{1},j_{2}\in K_{i}$ and for any $i\in[m]$. According to that result, the equation (40) can be rewritten as follows:
$\sum_{i=1}^{m}\sum_{j_{1}\in K_{i}}\exp{(b_{j_{1}})}\exp{((BW_{2}W_{1j_{1}}{\bm{X}})^{\top}{\bm{X}})}CW_{2}W_{1j_{1}}{\bm{X}}$
$=\sum_{i=1}^{m}\sum_{j_{2}\in K_{i}}\exp{(b_{*,j_{2}})}\exp{((BW_{*,2}W_{*,1j_{2}}{\bm{X}})^{\top}{\bm{X}})}CW_{*,2}W_{*,1j_{2}}{\bm{X}},$
for almost surely ${\bm{X}}$. That equation proves that
$\{W_{2}W_{1j_{1}}{\bm{X}}:j_{1}\in K_{i}\}=\{W_{*,2}W_{*,1j_{2}}{\bm{X}}:j_{2}\in K_{i}\},$
for any $i\in[m]$. That result indicates that
$\{W_{2}W_{1j_{1}}:j_{1}\in K_{i}\}=\{W_{*,2}W_{*,1j_{2}}:j_{2}\in K_{i}\}.$
As a consequence, we arrive at the following result:
$\sum_{i=1}^{m}\sum_{j_{1}\in K_{i}}\exp{(b_{j_{1}})}\delta_{W_{2}W_{1j_{1}}}=\sum_{i=1}^{m}\sum_{j_{2}\in K_{i}}\exp{(b_{*,j_{2}})}\delta_{W_{*,2}W_{*,1j_{2}}}.$
It is equivalent to $G\equiv\widetilde{G}_{*}$. As a consequence, we obtain the conclusion of the identifiability property of the function $f_{G}$. ∎
@@ Appendix C Related Work
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, [23] focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, [49, 46] established convergence rates for both density and parameter estimation in Softmax gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, [47, 48] employed least squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.